The no. of real solutions of x^4 – 5x^2 – 6x – 5= 0 os 
a. 0
​b. 1
​c. 3
​d. 4

Grade:11

1 Answers

Sher Mohammad IIT Delhi
9 years ago
(x^2-2)^2-x^2-6x-9=0
(x^2-2)^2-(x+3)^2=0
the function will be zero when
(x^2-2)^2=(x+3)^2
x^2-2=x+3 orx^2-2+x+3=0
x^2-x-5=0 0r x^2+x+1=0
the equation has two real solutions.
sher mohammad
faculty askiitians
b.tech, iit delhi